The vapor pressure of ethanol, C2H5OH, at 20 oC is 400 mmHg. Suppose that 6.5 grams of it are sealed in a 15.0-liter flask. Determine the mass of ethanol that enters the gas phase and the amount that remains in the liquid state upon equilibrium being established.
Answer #1

Let n moles of Ethanol evaporate

So, Using Ideal gas Law

PV = nRT

P= 400mmHg = 400/760 atm = 0.526 atm

T = 20C = 293K

R = 0.0821 atmL/mol.K

V = 15L

n = PV/RT

=( 0.526*15)/(0.0821*293)

=0.328 moles

Molar mass of ethanol = 46g/mol

So mass of Ethanol evaporated = 46*0.328 = 15.08 g

But we have only 6.5g

So this means that all the ethanol will vaporize to gas phase.